How many quarts of a 10% solution of basic acid must be added to 2 quarts of a 3% solution of basic acid to obtain a 4% solution?
What are the steps to answer this question. Thank you.

Answer by mananth(16949)

-------------- percent ---------------- Amount
Acid I 3 ---------------- 2 quarts
Acid II 10 ---------------- x quarts
Mixture 4 ---------------- 2 + x quarts

3 * 2 + 10 x = 4 ( 2 + x )
6 + 10 x = 8 + 4 x
10 x -4 x = -6 + 8
6 x = 2
/ 6
x = 0.33 quarts Acid II

m.ananth@hotmail.ca